OpenStudy (anonymous):

(t - 7) 2 = (t + 7) 2 + 168

OpenStudy (blurbendy):

(t-7)^2 = (t+y)^2 + 168?

OpenStudy (blurbendy):

or (t-7)*2 = (t+y)*2 + 168?

OpenStudy (anonymous):

(t-7)^2=(t+7)^2+168